When one side of a square is double and the other side is reduced by 2 units,the area of the resulting rectangle is 126 unit2 .Find the area, in unit2,of the square..

1
Expert's answer
2021-05-10T13:08:44-0400

Let x be the side of the square.

"2x(x-2)=126."

"x-2x-63=0."

"x=\frac{2+\sqrt{4+4*63}}{2}=9."

Area of the square: "A=x^2=9^2=81\;unit^2."